OpenStudy (anonymous):

find point P on the graph of the equation y=sqrt(x) closest to the point (5,0)

OpenStudy (anonymous):

i got 2/9 i dont know where i made a mistake

OpenStudy (anonymous):

a point on the curve looks like \((x,\sqrt{x})\) the square of the distance to (5,0) will be \[(x-5)^2+\sqrt{x}^2=x^2-10x+25+x\] \[=x^2-9x+35\] find the minimum values of this quadratic

OpenStudy (anonymous):

sorry that lasts line should have been \[x^2-9x+25\]

OpenStudy (anonymous):

i see where my mistake is

OpenStudy (anonymous):

good from there?

OpenStudy (anonymous):

just solve the quadratic and plug in the values to see which one is the minimum right?

OpenStudy (anonymous):

minimum is at the vertex, find it via \(-\frac{b}{2a}\)

OpenStudy (anonymous):

that is why the answer is \(\frac{9}{2}\)